[Yehonathan, Shruti]
First we turned on the relevant instruments for this experiment after the power shutdown:
- Main laser drivers and doubling cavity controller. We set the current to 2 A as we had it before.
- The waveguide TEC. We tried setting it to 60.99 C (for maximum efficiency) but the temperature ramps up much faster and over shoots the setpoint. So we had to do what we did earlier which was to adiabatically change the setpoint from room temperature and finally set it to something like 63 C so the actual measured temperature stabilizes at ~60.9 C. How do we change the PID parameters on this controller? The settings don't seem to allow for it.
- PD power supply, oscilloscopes, function generator, SR 560s lying nearby
Then we tried to probe further what was going on with the PDs (TL;DR not much made sense or was reproducible):
- Initially we were sending in 12 V from Ch1 of the Tenma power supply and
15 V to the mini-circuits RF amplifiers from Ch2 of the same Tenma. When we tried to observe the DC voltage levels (before amplification) and the noise (after amplification) but they did not make sense to us as described in the previous elog.
- Then we disconnected the RF amplifiers and switched the power supply for the PD transimpedance amps to Ch2 of the Tenma. Briefly we were able to see 2 V DC at each output (as expected for the ~1.3 mW of light incident on each PD) when measured with the multimeter. On the scope we were seeing that one of the PDs was dominated by 60 Hz fluctuations with a much lower
- We tried to switch it back to Ch1 but even the DC levels seen by the multimeter were bogus (~1 V for one of the channels and 2 V for the other)
- When we switched the power supply to a HP one taken from CTN, the DC levels on both channels seemed ~1V without any light and the noise somehow seemed lower with some incident light
Possible next steps
- Even though the AD 829 data sheet says it can be operated at up to 15 V, it is possible that we damaged both of them somehow when cycling the power. Elog 2324 also says that it was designed to be powered at
5 V. We could replace both op-amps and measure teh transimpedance TFs before and after the change
- Use different PDs. Maybe temporarily try with just the ThorLabs PDA20CS or similar with ~20 mW of LO power and measure the shot noise (possibly also squeezing).
|